Prove that the height of an AVL tree is bound by O(logn).

and What is the least number of nodes in an AVL tree of height 25? (all work on both questions provided)
Answer #1

`Hey,

Note: Brother if you have any queries related the answer please do comment. I would be very happy to resolve all your queries.

Proof: Let n(h) be the minimum number of internal nodes in an AVL tree of height h.

n(1)=1 and n(2)=2. So the lemma holds for h=1 and h=2.

Here comes the key point.

Consider an AVL tree of height h≥3 and the minimum number of nodes. This tree is composed of a root, and two subtrees. Since the whole tree has the minimum number of nodes for its height so do the subtrees. For the big tree to be of height h, one of the subtrees must be of height h-1. To get the minimum number of nodes the other subtree is of height h-2.
Why can't the other subtree be of height h-3 or h-4?
The height of siblings can differ by at most 1!

What the last paragraph says in symbols is that for h≥3,
n(h) = 1+n(h-1)+n(h-2)

The rest is just algebra, i.e. has nothing to do with trees, heights, searches, siblings, etc.

n(h) > n(h-1) so n(h-1) > n(h-2). Hence

n(h) > n(h-1)+n(h-2) > 2n(h-2)

Really we could stop here. We have shown that n(h) at least doubles when h goes up by 2. This says that n(h) is exponential in h and hence h is logarithmic in n. But we will proceed slower. Applying the last formula i times we get

For any i>0, n(h) > 2in(h-2i)    (*)

Let's find an i so that h-2i is guaranteed to be 1 or 2, since this would guarantee that n(h-2i) ≥ 1.
I claim i = ⌈h/2⌉-1 works.

If h is even h-2i = h-(h-2) = 2

If h is odd h-2i = h - (2⌈h/2⌉-2) = h - ((h+1)-2) = 1

Now we plug this value of i into equation (*) and get for h≥3

n(h) > 2in(h-2i)
   = 2⌈h/2⌉-1n(h-2i)
   ≥ 2⌈h/2⌉-1(1)
   ≥ 2(h/2)-1

Theorem: the height of an AVL tree storing n items is O(log(n)).

Proof: From the lemma we have n(h) > 2(h/2)-1.

Taking logs gives log(n(h)) > (h/2)-1 or

h < 2log(n(h))+2

Since n(h) is the smallest number of nodes possible for an AVL tree of height h, we see that h < 2 log(n) for any AVL tree of height h.

2)

For h=25

n>2^((h-2)/2)

So,

n=2897

Kindly revert for any queries

Thanks.